# Do Vision & Language Decoders use Images and Text equally? How Self-consistent are their Explanations?

### v2

@@ 4.1 Models and Data
As open access VLMs we use BakLLaVA, LLaVA-NeXT-Mistral, and LLaVA-NeXT-Vicuna [^3] We list exact links to the models in App A.1.:
BakLLaVA is a Mistral-7b-base Jiang et al. (2023) LLM augmented for VL processing with the LLaVA 1.5 Liu et al. (2023) architecture, which in turn builds on LLaVA Liu et al. (2024b) with two modifications: with i) a better visual encoder backbone, and with ii) training on academic VQA data with simple response formatting prompts.
LLaVA-NeXT-Mistral Liu et al. (2024a) in its Mistral-7b-base version improves LLaVa-1.5 by “increasing the input image resolution and training on an improved visual instruction tuning dataset to improve OCR and common sense reasoning”.
LLaVA-NeXT-Vicuna is the same as LLaVA-NeXT-Mistral, but with a Vicuna-7b Zheng et al. (2024) LLM. Vicuna is a LLaMA 1 (Touvron et al., 2023) finetuned on high-quality conversations.
@@
For multiple-choice tasks, we distinguish an easier and a harder setting: In pairwise multiple-choice, we prompt models to choose between captions and unfitting captions (prompt in App. A.2). In image-sentence alignment multiple-choice, we prompt them to say whether an image and a sentence match or mismatch (prompt in App. A.2).
@@ 4.2 Metrics
Following Parcalabescu et al. (2022), we use four performance metrics: caption precision ($p_{c}$) measuring how well models identify the correct examples in image-sentence alignment multiple-choice with just captions; foil precision ($p_{f}$) measuring how many of the foiled cases are correctly identified in image-sentence alignment multiple-choice; overall accuracy ($acc$) on all classes (foil and correct in the image-sentence alignment multiple-choice setting [^4] Because all data we test on contains 50% matching and 50% mismatching pairs, $acc$ is the average of $p_{c}$ and $p_{f}$.); pairwise accuracy ($acc_{r}$), measuring how often the model chooses the caption over the foil in pairwise multiple-choice. The pairwise accuracy $acc_{r}$ is more permissive than $acc$, because for $acc_{r}$, the model has both the caption and the foil in its input for an image. So, the model can directly compare the two, and exploit linguistic differences between them. Encoders, by construction, do not accept both caption and foil next to the image input, and therefore Parcalabescu et al. (2022) computed $acc_{r}$ by checking whether the image-sentence alignment score is greater for a correct image-text pair than for its foil counterpart.
@@
We assess the self-consistency of VLMs with CC-SHAP in both post-hoc and CoT explanation settings. CC-SHAP can be directly applied to VL decoders, provided that Shapley values can be calculated – for which we write the necessary code. CC-SHAP is a continuous value between -1 (opposite self-consistency) and 1 (perfect self-consistency). 0 is no self-consistency.
Additionally, we implement six existing (edit-based) tests for VLMs: Counterfactual Edits, Biasing Features, and Corrupting CoT: Adding Mistakes, Early Answering, Filler Tokens, and Paraphrasing. We report the percentage of samples deemed to be self-consistent by these tests.
@@
We also compute MM-SHAP for the VLMs in prediction and explanation settings (both post-hoc and CoT). Unless specified otherwise, we conduct the MM-SHAP and self-consistency experiments on 100 random samples from each dataset (and instrument of VALSE), due to computational demands outlined in §6.1. We provide variance estimations for our results in App. A.3, Figures 7 and 8. Performance results on VALSE ($acc$, $p_{c}$, $p_{f}$, $acc_{r}$ in the next section §4.3) are computed on all samples from VALSE (not a subset).
@@ 4.3 VALSE Results with VLM Decoders
Results for all VL decoders on all samples from VALSE are listed in Tab. 1. We compare in Fig. 1 to prior results on VALSE with VL encoders as published by Parcalabescu et al. (2022).
@@
– Tab. 1. Best average zero-shot results according to $acc_{r}$ are achieved by LLaVA-NeXT-Mistral (85%), followed by BakLLaVA (83%) and LLaVA-NeXT-Vicuna (79%). With $acc_{r}$, the results are generally very strong, except for the counting adversarial instrument (numbers underlined in Tab. 1). This underscores that VL decoders are using linguistic priors, such as with the counting adversarial test where the captions contain small numbers, while the foils contain large numbers – test designed to counteract VLMs that are biased towards small numbers, which are more frequent in training data.
Results for $acc$ are typically near-random. However, object-centred instruments such as existence and Foil-it yield somewhat better outcomes. The significant difference between overall higher $acc_{r}$ and lower $acc$ results suggests that VL decoders rely on linguistic priors to solve VALSE. The fanned-out $p_{c}$ and $p_{f}$ metrics show that LLaVA-NeXT-Vicuna is biased towards predicting that a sentence is a correct description of the image. BakLLaVA and LLaVA-NeXT-Mistral are biased towards predicting the opposite and better identify foils. Both models contain a Mistral-7B LLM, which explains why they share a tendency.
Fig. 1 shows that the decoder models of 2024 are performing better than the encoder models of 2019-2021, but only in terms of $acc_{r}$, where decoders must choose between the caption and the foil in the pairwise multiple-choice setting and can exploit linguistic and plausibility biases by directly comparing caption and foil. Importantly, judging by the more challenging metric $acc$, decoders do not generally outperform encoder models.
@@ 4.4 Evaluating the Multimodal Degree and Self-Consistency of VLMs
We visualise key metrics: Fig. 4 shows accuracy and MM-SHAP scores for VLMs on VALSE, VQA, MSCOCO, GQA, and GQA balanced. Fig. 5 shows CC-SHAP post-hoc and CC-SHAP CoT scores on VALSE, as well as Counterfactual Edits test results. Full results for all VLMs and tests, applied to generation tasks (VQA, GQA, GQA balanced) and the MSCOCO image-sentence alignment task are listed in App. A.3 Tab. 2. Tab. 3 shows the results for the multiple-choice tasks of VALSE and FoilIt.
@@
VL decoders in Fig. 4 show pronounced reliance on the text modality when generating answers. Specifically, BakLLaVA, LLaVA-NeXT-Mistral and LLaVA-NeXT-Vicuna exhibit text degrees (T-SHAP prediction) of 87%, 97% and 89% respectively in VQA, with similar trends on GQA. However, the text usage on GQA balanced even increases, with BakLLaVA, LLaVA-NeXT-Mistral and LLaVA-NeXT-Vicuna having a T-SHAP prediction of 90%, 96% and 90% respectively. This increase is in accordance with known stronger linguistic biases in GQA compared to GQA balanced.
@@
As seen in the previous section from Fig. 4, in VL decoders, the text modality predominates during prediction generation, with all T-SHAP prediction values at 89% and higher. From the explanation setting of our experiments, we gain more insight: the image modality becomes more influential during explanation generation compared to prediction, leading to a notable decrease in text modality contributions by 4 to 30 percentage points [^5] This is not due to the longer text generation, because the method from §3 ensures similar sequence lengths between text inputs and image patches in interpretation.. The decrease is even larger in CoT than in post-hoc explanation. Also, VLMs perform poorly in making predictions with CoT, indicating lesser CoT abilities relative to LLMs. This is likely due to the multimodal training data being less challenging, less linguistically diverse and lacking details, compared to language-only corpora (McKinzie et al., 2024). Despite this, MM-SHAP – which complements accuracy and works directly with probabilities (Lyu et al., 2024) – effectively assesses their multimodal capacity even under low accuracy conditions.
@@
Fig. 5 (a) and (b) (and Tab. 3 with full results) show varying scores across VALSE instruments (multiple-choice), with most CC-SHAP scores being negative. This indicates a misalignment between the contributions when VLMs predict and explain, suggesting that VLMs are less self-consistent than the LLMs studied by Parcalabescu and Frank (2024). This aligns with observations that the image is used more in explanation generation than in prediction, indicating a model inconsistency: why can it give answers without relying much on the image, yet turn to the image to explain its already-made decisions? This is also the case on individual examples, cf. Tab. 6.
In experiments on VQA, GQA, and GQA balanced (cf. Tab. 2), BakLLaVA and LLaVA-NeXT-Vicuna exhibit positive CC-SHAP in generative tasks. This is highlighted in green in Tab. 2, and is also attested in individual instances, e.g., in Tab. 4 for BakLLaVA – contrasting with their negative scores in multiple-choice from Fig. 5 (a) and (b), and Tab. 3 (cf. Tab. 8 for an individual sample).
@@
CC-SHAP post-hoc scores are closer to zero for simpler VALSE instruments, such as noun phrases, counting small numbers, and existence – although LLaVA-NeXT-Vicuna is an exception with a notable negative score of $-0.08$. In contrast, tasks designed to be challenging, such as counting adversarial, counting balanced, and coreference-clean, yield lower CC-SHAP post-hoc scores. Also, instruments with high plausibility bias – spatial relations, action replacement, and actant swap – show more negative scores, with LLaVA-NeXT-Mistral having the most negative ones. This is consistent with our observations with T-SHAP, namely that models primarily rely on text for predictions but shift their focus to the image for explanations, as there are high T-SHAP prediction values for action replacement and actant swap – blue in Tab. 3.
@@
Results from edit-based tests on generative tasks (including Biasing Features, Adding Mistakes, Early Answering, Filler Tokens, and Paraphrasing), in Tab. 2, show extreme self-consistency scores – either 0% or 100%. These tests modify model inputs and examine whether the output changes, which is easy to check in multiple-choice tasks, therefore the edit-based tests deliver meaningful results in Tab. 2 for MSCOCO, or in Tab. 3 for VALSE. However, generative tasks require semantic evaluation to determine if the output remains consistent, a process that is complex because the amount of tolerable output variation is sample-dependent. For instance, in the VQA sample from Tab. 5, BakLLaVA outputs that the horse is “on the sidewalk”, and post-edit emits “city intersection”, which is in fact more accurate. Humans cannot judge whether the model actually meant the same thing and whether the model was self-consistent or not – after all, why did the insertion of “trial and error” improve its answer? We do not know and its inner workings remain opaque. Cf. App. A.4 for more instances from these tests (including CC-SHAP) on actual samples.
@@
In App. A.4 (Tables 4 to 11), we provide examples of model generated explanations. There is no human annotation and ground truth for explanation faithfulness and model self-consistency. Neither previous work interested in faithfulness, nor we, evaluated the plausibility of generated explanations, because plausibility and faithfulness are orthogonal (Jacovi and Goldberg, 2020). Some prior tests do not even examine the explanation (such as the Corrupting CoT tests), others only search for specific keywords in them (e.g., the Counterfactual Edit). While not being able to judge the plausibility of the explanations at content level, our method, however, takes into account how much input tokens contribute in generating it, and compares this to the input tokens’ contributions when making the prediction. By inspecting the examples from App. A.4 with CC-SHAP, we can see whether the model really uses the image regions and text tokens corresponding to the concepts it mentions in the explanations. Yet, it is not possible to say exactly how positive or negative these contributions should be (although certainly non-zero), as self-consistency measures such as CC-SHAP do not reach too deeply into a model’s internals.
@@ 5 Conclusions and Future Work
Several open questions remain, e.g., why VL decoders are predominantly text-centred. Is this a result of their training and architecture, or does it stem from task instances containing excessive linguistic cues? Future work could clarify this by designing datasets devoid of plausibility biases and other linguistic indicators. Another point is the lower self-consistency of VL decoders compared to LLMs. Studies could investigate the models’ internals to determine whether this indicates a lack of self-consistency or an exploitation of data biases.
6 Limitations
This work focuses on assessing the performance of VLMs designed for English. Also, it measures the self-consistency of natural language explanations that are generated by such models. The following limitations can be relevant for future work.
6.1 Compute Requirements
Future work might be interested in reducing the compute requirements solicited by self-consistency tests. CC-SHAP needs around 7 minutes to run an example on a NVIDIA A40 48 GB GPU, and the other tests need around 1.5 minutes. CC-SHAP requires 25 minutes to compute self-consistency per example on a NVIDIA A40 48 GB GPU for LLaVA-NeXT models running with FlashAttention-2 (Dao, 2023) and quantisation (Dettmers et al., 2023) (otherwise the memory requirements exceed 48 GB), while the other tests need 7 minutes. VLMs have larger sequence length than LLMs, because they need to process many image tokens. LLaVA-NeXT variants even more so, because they process the image at five different resolutions (increasing the image sequence length by a factor of five). CC-SHAP needs to run multiple model inferences to estimate Shapley values, while the other self-consistency tests require at most two model inferences (e.g., Biasing Features Turpin et al. (2023)). We think that CC-SHAP’s compute time is well invested, since it is a more effective measure: it adds interpretability aspect and delivers better scores in settings such as VQA, where other tests deliver either 0% or 100% (cf. Tab. 2).
6.2 Standard Deviation of our Results
We provide variance estimations for our performance and self-consistency results on a representative subset of our experiments in App. A.3.
6.3 Further Investigations on Methodology
While the normalisation step described in Section 3 is needed to compare values that otherwise would have different ranges, the question of aggregation is more open. Averaging is a good strategy because we expect all tokens to be equally important. At present, we adopted averaging as a strategy, without being able to certify its superiority relative to other options. In our experiments it seemed to work out rather well. However, future work may consider and investigate other options.
6.4 Closed Source Models

### v4

@@ 4.1 Models and Data
We use four VLMs for our experiments, each consisting of 7 billion parameters: BakLLaVA, LLaVA-NeXT-Mistral, LLaVA-NeXT-Vicuna, and mPLUG-Owl3 – Links to the models are in Appendix B.3:
  • BakLLaVA – appeared October 2023 – is a Mistral-7b-base (Jiang et al., 2023) LLM augmented for VL processing with the LLaVA 1.5 (Liu et al., 2023) architecture, which in turn builds on LLaVA (Liu et al., 2024b) with two modifications: with i) a better visual encoder backbone, and with ii) training on academic VQA data with simple response formatting prompts.
  • LLaVA-NeXT-Mistral – appeared January 2024 (Liu et al., 2024a) – in its Mistral-7b-base version improves LLaVa-1.5 by “increasing the input image resolution and training on an improved visual instruction tuning dataset to improve OCR and common sense reasoning”.
  • LLaVA-NeXT-Vicuna is the same as LLaVA-NeXT-Mistral, but with a Vicuna-7b (Zheng et al., 2024) LLM. Vicuna is a LLaMA 1 (Touvron et al., 2023) finetuned on high-quality conversations.
  • mPLUG-Owl3 – appeared August 2024 (Ye et al., 2024) is the most recently published model and was not included in the first version of this work. It marks the beginning of a new generation of VLMs, being trained on lengthy video and not just (interleaved) image and text data.
@@
In multiple-choice, we define an easier and a harder setting: In pairwise multiple-choice, prompted models choose between captions and unfitting captions. In image-sentence alignment multiple-choice, we prompt them to say whether an image and a sentence match or not (prompts in B.4).
@@ 4.2 Metrics
Following Parcalabescu et al. (2022), we use four performance metrics:
  • Caption precision ($p_{c}$) measuring how well models identify the correct examples in image-sentence alignment multiple-choice with just captions;
  • Foil precision ($p_{f}$) measuring how many of the foiled cases are correctly identified in image-sentence alignment multiple-choice;
  • Overall accuracy ($acc$) on all classes (foil and correct in the image-sentence alignment multiple-choice setting [^3] All data we test on contains 50% matching and 50% mismatching pairs, so $acc$ is the average of $p_{c}$ and $p_{f}$.);
  • Pairwise accuracy ($acc_{r}$), measuring how often the model chooses the caption over the foil in pairwise multiple-choice setting.
The pairwise accuracy $acc_{r}$ setting is more permissive than $acc$, because for $acc_{r}$, the model has both the caption and the foil in its input for an image. So, the model can directly compare the two, and exploit linguistic differences between them. Encoders, by construction, do not accept both caption and foil next to the image input, and therefore Parcalabescu et al. (2022) computed $acc_{r}$ by checking whether the image-sentence alignment score is greater for a correct image-text pair than for its foil counterpart.
@@
We assess the self-consistency of VLMs with CC-SHAP in both post-hoc and CoT explanation settings. CC-SHAP as presented by (Parcalabescu & Frank, 2024) can be directly applied to VL decoders, provided that Shapley values can be calculated – for which we produced the necessary code. CC-SHAP is a continuous value between -1 (opposite self-consistency) and 1 (perfect self-consistency). 0 is no self-consistency. See Appendix B.2 for a recap of CC-SHAP.
Additionally, we implement six existing (edit-based) tests for the VLMs [^4] Due to time constraints and significant differences in mPLUG-Owl3’s code functionality, we did not implement these tests for mPLUG-Owl3 and instead rely solely on CC-SHAP for self-consistency evaluations.: Counterfactual Edits, Biasing Features, and Corrupting CoT: Adding Mistakes, Early Answering, Filler Tokens, and Paraphrasing. We report the percentage of samples deemed to be self-consistent by these tests.
@@
We also compute MM-SHAP for the VLMs in answer and explanation settings (both post-hoc and CoT). Unless specified otherwise, we conduct the MM-SHAP and self-consistency experiments on 100 random samples from each dataset (and VALSE instruments) due to computational demands outlined in §A.2. We provide variance estimations for our results in § B.5, Fig. 7 and 8. Performance results on VALSE ($acc$, $p_{c}$, $p_{f}$, $acc_{r}$ in §4.3) are computed on all samples from VALSE (not a subset).
@@ 4.3 VALSE Results with VLM Decoders
Results for all VL decoders on all VALSE samples are in Tab. 1. We compare in Fig. 1 to average results over all phenomena in VALSE with VL encoders as published by Parcalabescu et al. (2022).
@@
– Tab. 1. Best average zero-shot results according to $acc_{r}$ are achieved by mPLUG-Owl3 (88%), followed by LLaVA-NeXT-Mistral (85%), BakLLaVA (83%), and LLaVA-NeXT-Vicuna (79%). With $acc_{r}$, the results are generally very strong for all models, except on the counting adversarial instrument (numbers underlined in Tab. 1). This underscores that VL decoders (except mPLUG-Owl3 which is the most recent one) are using linguistic priors, such as with the counting adversarial test where the captions contain small numbers, while the foils contain large numbers – test designed to counteract VLMs that are biased towards small numbers, which are more frequent in training data.
Results for $acc$ are typically near-random for all models except mPLUG-Owl3. However, object-centred instruments, such as existence and Foil-it, yield better outcomes. The large difference between overall higher $acc_{r}$ and lower $acc$ results suggests that VL decoders rely on linguistic priors to solve VALSE. The fanned-out $p_{c}$ and $p_{f}$ metrics show that LLaVA-NeXT-Vicuna is biased towards predicting that a sentence is a correct description of the image. BakLLaVA and LLaVA-NeXT-Mistral rather predict the opposite and better identify foils. Both models contain a Mistral-7B LLM, which explains why they share a tendency. mPLUG-Owl3, however, does not exhibit a strong bias towards either $p_{c}$ or $p_{f}$. While it generally performs better on $p_{c}$ by favoring the interpretation that a sentence accurately describes the image, it sometimes excels on $p_{f}$, as seen with tasks like counting. In VALSE’s adversarial settings, such as counting adversarial, mPLUG-Owl3 still struggles, though its difficulties appear to stem from challenges in holistic image-text understanding rather than linguistic biases, as its $acc_{r}$ remains unaffected in these cases.
Fig. 1 shows that the decoder models of 2024 are performing better than the encoder models of 2019-2021 in $acc_{r}$, where decoders must choose between the caption and the foil in the pairwise multiple-choice setting and can exploit linguistic and plausibility biases by directly comparing caption and foil. Importantly, judging by the more challenging metric $acc$, decoders do not generally outperform encoder models.
@@ 4.4 Evaluating the Multimodal Degree and Self-Consistency of VLMs
We visualise key metrics: Fig. 4 shows accuracy and MM-SHAP scores for VLMs on VALSE, VQA, MSCOCO, GQA, and GQA balanced. Fig. 5 shows CC-SHAP post-hoc and CC-SHAP CoT scores on VALSE, as well as Counterfactual Edits test results. Full results for all VLMs and tests, applied to generation tasks (VQA, GQA, GQA balanced) and the MSCOCO image-sentence alignment task are listed in App. B.5 Tab. 2. Tab. 3 shows the results for the multiple-choice tasks of VALSE and FoilIt.
@@
The tested VL decoders in Fig. 4 show pronounced reliance on the text modality when generating answers. Specifically, BakLLaVA, LLaVA-NeXT-Mistral, LLaVA-NeXT-Vicuna, and mPLUG-Owl3 exhibit text degrees (T-SHAP answer) of 87%, 97%, 89%, and 85% respectively in VQA, with similar trends on GQA. However, the text usage on GQA balanced even increases, with BakLLaVA, LLaVA-NeXT-Mistral and LLaVA-NeXT-Vicuna having a T-SHAP answer of 90%, 96% and 90% respectively. This increase is in accordance with known stronger linguistic biases in GQA compared to GQA balanced.
@@
As seen in the previous section from Fig. 4, in VL decoders, the text modality predominates during answer generation, with all T-SHAP answer values at 85% and higher. From the explanation setting of our experiments, we gain more insight: the image modality becomes more influential during explanation generation compared to answer, leading to a statistically significant decreases in text modality contributions by 1 to 30 percentage points [^5] This is not due to the longer text generation, because the method from §3 ensures similar sequence lengths between text inputs and image patches in interpretation.. The decrease is even larger in CoT than in post-hoc explanation. Also, all tested VLMs perform significantly worse when answering with CoT, except for mPLUG-Owl3, which shows only a slight decline. This suggests that VLMs have weaker CoT capabilities compared to LLMs. mPLUG-Owl3 appears to benefit from its broader training data, which includes videos. For the other models, this limitation is likely due to their multimodal training data being less challenging, less linguistically diverse, and lacking the detail found in language-only corpora (McKinzie et al., 2024). Despite low accuracies, MM-SHAP – which complements accuracy and works directly with probabilities (Lyu et al., 2024) – effectively assesses their multimodal capacity even under low accuracy conditions.
@@
Fig. 5 (a, b) (and Tab. 3 with full results) show varying scores across VALSE instruments (multiple-choice), with BakLLaVA, LLaVA-NeXT-Vicuna and -Mistral showing negative CC-SHAP. This indicates misalignment between the contributions when VLMs predict and explain, suggesting that some VLMs are less self-consistent than the LLMs tested by Parcalabescu & Frank (2024). This is partly explained by these models’ relying more on image information when generating explanations than answers (cf. Fig. 4), indicating a model inconsistency: why can models give answers without relying much on the image, yet turn to the image to explain its already-made decisions? This is also the case on individual examples, cf. Tab. 6. In contrast, mPLUG-Owl3 shows a smaller gap in image usage between explanation and answer generation, enabling it to achieve positive CC-SHAP—assuming convergent behavior between answers and explanations beyond modality focus. It stands out as a superior model, exemplifying the advancements as of August 2024, surpassing its LLaVA-based counterparts (the latest from January 2024) in performance and self-consistency.
In experiments on generative tasks (VQA, GQA, and GQA balanced), mPLUG-Owl3 continues showing positive CC-SHAP (cf. Tab. 2) – BakLLaVA and LLaVA-NeXT-Vicuna do too. This is highlighted in green in Tab. 2, and is also attested in individual instances, e.g., in Tab. 4 for BakLLaVA – contrasting with their negative scores in multiple-choice from Fig. 5 (a) and (b), and Tab. 3 (cf. Tab. 8 for an individual sample).
@@
Negative CC-SHAP post-hoc scores are larger for simpler VALSE instruments, such as noun phrases, counting small numbers, and existence – although LLaVA-NeXT-Vicuna is an exception with a notable negative score of $-0.08$. In contrast, tasks designed to be challenging, such as counting adversarial, counting balanced, and coreference-clean, yield lower CC-SHAP post-hoc scores. Also, instruments with high plausibility bias – spatial relations, action replacement, and actant swap – show more negative scores, with LLaVA-NeXT-Mistral having the most negative ones. This is consistent with our observations with T-SHAP, namely that models primarily rely on text for answers but shift their focus to the image for explanations, as there are high T-SHAP answer values for action replacement and actant swap – blue in Tab. 3.
@@
Results from edit-based tests on generative tasks (including Biasing Features, Adding Mistakes, Early Answering, Filler Tokens, and Paraphrasing), in Tab. 2, show extreme self-consistency scores – either 0% or 100%. This inconsistent scoring mirrors the findings of Parcalabescu & Frank (2024) for LLMs. As discussed in the related work section, these tests have several issues. One key limitation is their reliance on modifying model inputs and observing whether the output changes. While these tests modify model inputs and examine whether the output changes, which is easy to check in multiple-choice tasks, therefore the edit-based tests deliver meaningful results in Tab. 2 for MSCOCO, or in Tab. 3 for VALSE. However, generative tasks require semantic evaluation to determine if the output remains consistent, a process that is complex because the amount of tolerable output variation is sample-dependent. For instance, in the VQA sample from Tab. 5, BakLLaVA outputs that the horse is “on the sidewalk”, and post-edit emits “city intersection”, which is in fact more accurate. Humans cannot judge whether the model actually meant the same thing and whether the model was self-consistent or not – after all, why did the insertion of “trial and error” improve its answer? We do not know and its inner workings remain opaque. Cf. App. B.6 for more instances from these tests (including CC-SHAP) on actual samples.
@@
In App. B.6 (Tables 4 to 11), we provide examples of model generated explanations. There is no human annotation and ground truth for explanation faithfulness and model self-consistency. Neither previous work interested in faithfulness, nor we, evaluated the plausibility of generated explanations, because plausibility and faithfulness are orthogonal (Jacovi & Goldberg, 2020). Some prior tests do not even examine the explanation (such as the Corrupting CoT tests), others only search for specific keywords in them (e.g., the Counterfactual Edit). While not being able to judge the plausibility of the explanations at content level, our method, however, takes into account how much input tokens contribute in generating it, and compares this to the input tokens’ contributions when giving the answer. By inspecting the examples from App. B.6 with CC-SHAP, we can see whether the model really uses the image regions and text tokens corresponding to the concepts it mentions in the explanations. Yet, it is not possible to say exactly how positive or negative these contributions should be (although certainly non-zero), as self-consistency measures such as CC-SHAP do not reach too deeply into a model’s internals.
@@ 5 Conclusions and Future Work
Outlook. Several open questions remain, e.g., why VL decoders are predominantly text-centred. Is this a result of their training and architecture, or does it stem from task instances containing excessive linguistic cues? Future work could clarify this by designing datasets devoid of plausibility biases and other linguistic indicators. Another point is the lower self-consistency of VL decoders compared to LLMs. Studies could investigate the models’ internals to determine whether this indicates a lack of self-consistency or an exploitation of data biases.
Limitations. We assessed the performance and multimodal degree of VLMs and measured the self-consistency of natural language explanations generated by such models. Cf. App. A for a discussion of limitations, including model selection, compute requirements and method design choices.
Acknowledgements
We would like to thank the anonymous reviewers for their useful suggestions. Thanks go to Nils Trost for assisting with the visualisations. We acknowledge the support given by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG.
Appendix A Appendix A: Limitations
The following limitations can be relevant for future work.
A.1 Closed Source Models